# CRoP: Context-wise Robust Static Human-Sensing Personalization

### v4

@@ 6.1.1. Pre-Processing of the Datasets
6.1.2. Training of the Generic models
PERCEPT-R
: For the identified 16 participants for our personalization evaluation, their speech data were collected longitudinally, meaning their data could be separated into available and unseen contexts versus other speakers in the corpus who only had speech data available from a one-time point. The generic model is trained for the remaining 499 participants (having / r / sounds disorder) using person-disjoint validation and test sets. The aim of this dataset is to identify the correctness of / r / sounds.
Model: In line with the literature Benway et al. (2023), we tried several model architectures such as CNN, DNN, BILSTM, etc, whose number of parameters were identified using grid search. Among those, the biLSTM model containing 4 bidirectional LSTM layers followed by 5 linear layers, accompanied by a Hardswish activation layer, was identified as the one that exhibited the best results for the generic data and was used for this study.
WIDAR:
We chose 3 users for personalization since these were the only users whose data was collected in both rooms. Since the number of users in WIDAR is very small, the exclusion of all the 3 users for training the generic model would have resulted in substandard models. So, For each user, we generated different generic models by using data from the other 16 users with a 14/2 person disjoint random split for the train and validation set. Our classification target was the 6 gesture classes: 0,1,2,3,5 and 8, corresponding to push, sweep, clap, slide, draw a circle, and draw zigzag, respectively, as evaluated in the original work (Zhang et al., 2022).
Model: The model used for WIDAR follows the LeNet architecture (Zhang et al., 2022), which contains three 2D convolutional layers followed by two linear layers. Each of these layers, except the final classification layer, is followed by a ReLU activation layer.
ExtraSensory:
We chose 5 users for personalization, and the generic model is trained on 42 users, with 10 users being left out for person-disjoint validation. The two target classes are walking and sitting.
Model: The model follows a CNN-GRU-based architecture used in HAR literature (Gong et al., 2019; Hao et al., 2022; Shen et al., 2022; Yao et al., 2017). The model consists of three batch-normalized 1D convolution layers followed by a linear layer that feeds into a batch-normalized recursive (GRU) layer and two linear layers to generate embeddings. For the classification head, two linear layers were used.
Stress Sensing Dataset:
For this binary task, three users with data across all contexts were selected for personalization. The generic model was trained on data from 21 users, with six additional users for disjoint validation and test sets.
Model: The model uses a simple multi-layer-perceptron (MLP) architecture (Eren and Navruz, 2022; Dzieżyc et al., 2020; Wang et al., 2017) consisting of 3 linear layers with hidden size of 128.
@@ 6.2. Metrics for evaluation
In summary, the metric $\Delta_{P}$ suggests how well the model $\mathcal{M}^{P^{a}_{i}}_{\theta}$ performs as compared to the generic model $\mathcal{M}^{G}_{\theta}$. Since the generic model has not learned the person-specific patterns. This metric quantifies CRoP’s ability to learn person-specific patterns. On the other hand, conventionally finetuned models $\mathcal{M}^{C^{a}_{i}}_{\theta}$ may learn person-specific patterns and forget the generic information of different contexts. Thus, the metric $\Delta_{G}$ quantifies CRoP’s ability to retain generic information.
@@ 6.3. Comparison with State-of-the-art
To demonstrate the efficacy of CRoP in achieving personalization $\Delta_{P}$ while maintaining generalization $\Delta_{G}$, we compare CRoP with 5 state-of-the-art approaches SHOT (Liang et al., 2020), PackNet (Mallya and Lazebnik, 2018), Piggyback (Mallya et al., 2018), CoTTA (Wang et al., 2022b), and PTN (Burns et al., 2022).
@@
Figure 4. Detailed results for Stress sensing dataset User 3 under Scenario 1 using F1 score as the evaluation metric
@@ 6.3.3. Inferior generic model performance leads to higher gain in personalization
For ExtraSensory and Stress-sensing datasets, certain users experience sub-optimal performance with the generic model $\mathcal{M}^{G}_{\theta}$ and the personalized finetuning helped not only in available context but also in the unseen contexts. For instance, the generic model showed suboptimal performance for certain users such as user ‘80’ for ExtraSensory dataset under Scenario 1 and user ‘3’ for Stress-sensing dataset. Such users show higher benefit even in the unseen context when personalized using available context data, highlighting the importance of user-specific patterns. The details explanation of these results are provided in Appendix B.
@@ 7.1. How different steps of CRoP facilitate generalizability
This section empirically discusses how each step of CRoP (Algorithm 1) facilitates intra-user generalizability, signifying similarity in the model’s behavior towards available $\mathcal{C}_{a}$ (available during personalization finetuning) and unseen contexts $\mathcal{C}_{u}$.
Shi et al. (2021) introduced the use of gradient inner product (GIP) to estimate the similarity between a model’s behavior across different domains. If $G_{i}$ and $G_{j}$ represent the gradient incurred by the model for Domains $D_{i}$ and $D_{j}$, then the sign of the product $G_{i}*G_{j}$ represents whether the model treats two domains similarly or not. For instance, $G_{i}*G_{j}>0$ signifies that the gradient for both domains has the same direction. We used GIP to quantify generalization. A higher GIP value for a personalized model across available ($\mathcal{C}_{a}$) and unseen contexts ($\mathcal{C}_{u}$) indicates more similar behavior toward both domains, indicating higher intra-user generalizability. GIP is measured as: $\|\sum_{i}G_{i}\|^{2}-\sum_{i}\|G_{i}\|^{2}$.
Figure 5. Variation of GIP at different stages of CRoP signified by the lines in Algorithm 1
Figure 5 shows that fine-tuning the generic model (Algorithm 1 – Step $2$) on Context $\mathcal{C}_{a}$, optimizes the model for this context, leading to a negative GIP, indicating a greater discrepancy between two contexts. Since model pruning results in generalization (Jin et al., 2022), an increase in GIP value can be observed in the pruned model (Step $3$).
On further analysis, we found that the model complementary to the pruned model (that is, the parameters that were removed) also contributed towards inter-context behavior discrepancy (negative GIP value). However, the same parameters in the generic model (that are replaced in Step $4$) formed a more generalizable set of weights, i.e., GIP $\geq 0$. Thus, the model mixing step (Step $4$) introduces further generalizability (GIP $\geq 0$) in the personalized model.
Finally, as discussed in Section 5.2 and following Section 7.2, the objective of the final finetuning step is to recover the performance loss, and it doesn’t aim to enhance generalizability further, as also shown in Figure 5.
7.2. Justification for the requirement of final finetuning step (Algorithm 1 Line 5) for some users
Figure 6. Activation maps in $\mathcal{C}_{a}$ for Scenario 1
In Section 5.2, we discussed that the Mixed Model from Algorithm 1 Line 4 suffers accuracy loss due to altered activation paths, as it combines weights from the generic and personalized stages. To illustrate, we compare the performance and activation maps of the model states at step 2 ($\mathcal{M}^{P^{a}_{i}}{\theta^{\prime}}$) and the Mixed Model ($\mathcal{M}^{P^{a}_{i}}{\theta^{\prime\prime}}$).
Table 5 shows that the Mixed Model $\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime\prime}}$ does show an average improvement of $14.43\%$ in inference accuracy for the unseen context $\mathcal{C}_{u}$, due to retainment of generic model weights; however, there is significant loss in the available context $\mathcal{C}_{a}$ as compared to the finetuned model $\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}}$. We attribute this loss to both pruning-induced information loss and inconsistent changes in activation paths during model mixing. Figure 6 shows the change in the activation map for a sample from context $\mathcal{C}_{a}$, which was correctly classifier by $\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}}$; however, it got misclassified after the model mixing step. Significant differences in activation maps can be observed for this sample. This highlights the need for finetuning to recover lost information and restore consistent activation paths.
Table 5 further shows that the final finetuning step indeed helps recover the loss in the available context $\mathcal{C}_{a}$. Although this may slightly reduce accuracy in the unseen context due to overwriting some generic patterns, it provides an average improvement of 4.25 percentage points in the unseen context $\mathcal{C}_{u}$ as compared to the model achieved after first finetuning (Algorithm 1 Line 2). This shows that the final finetuning is necessary for some users to regain performance lost in available context $\mathcal{C}_{a}$.
Model $\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}}$ $\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime\prime}}$ $\mathcal{A}(\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime\prime}},\mathcal{C})-\mathcal{A}(\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}},\mathcal{C})$ $\mathcal{M}^{P^{a}_{i}}_{\theta}$ $\mathcal{A}(\mathcal{M}^{P^{a}_{i}}_{\theta},\mathcal{C})-\mathcal{A}(\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}},\mathcal{C})$
User $\mathcal{C}_{a}$ $\mathcal{C}_{u}$ $\mathcal{C}_{a}$ $\mathcal{C}_{u}$ $\mathcal{C}_{a}$ $\mathcal{C}_{u}$ $\mathcal{C}_{a}$ $\mathcal{C}_{u}$ $\mathcal{C}_{a}$ $\mathcal{C}_{u}$
0 87.06 65.02 70.93 74.64 -16.13 +9.62 83.67 69.53 -3.39 +4.51
1 89.38 44.38 74.23 72.70 -15.15 +28.32 86.41 54.45 -2.97 +10.07
2 71.88 64.45 60.09 69.79 -11.79 +5.34 77.02 62.63 +5.14 -1.82
Average -14.36 +14.43 -0.41 +4.25
Table 5. Performance Comparison of model states after initial finetuning (Algorithm 1 line 2), model mixing (Algorithm 1 line 4) and final finetuning (Algorithm 1 line 5) for WIDAR dataset under Scenario 1

### v5

@@ 6.1.1. Pre-Processing of the Datasets
6.1.2. Data segmentation for Personalization and Generic Model Training
PERCEPT-R: For the identified 16 participants for our personalization evaluation, their speech data were collected longitudinally, meaning their data could be separated into available and unseen contexts versus other speakers in the corpus who only had speech data available from a one-time point. The generic model is trained for the remaining 499 participants (having / r / sounds disorder) using person-disjoint validation and test sets. The aim of this dataset is to identify the correctness of / r / sounds.
WIDAR: We chose 3 users for personalization since these were the only users whose data was collected in both rooms. Since the number of users in WIDAR is very small, the exclusion of all the 3 users for training the generic model would have resulted in substandard models. So, For each user, we generated different generic models by using data from the other 16 users with a 14/2 person disjoint random split for the train and validation set. Our classification target was the 6 gesture classes: 0,1,2,3,5 and 8, corresponding to push, sweep, clap, slide, draw a circle, and draw zigzag, respectively, as evaluated in the original work (Zhang et al., 2022).
ExtraSensory: We chose 5 users for personalization, and the generic model is trained on 42 users, with 10 users being left out for person-disjoint validation. The two target classes are walking and sitting.
Stress Sensing: For this binary task, three users with data across all contexts were selected for personalization. The generic model was trained on data from 21 users, with six users for disjoint validation and test sets.
Generic model training: We employ the model architectures and training methods used in the literature corresponding to each dataset: PERCEPT-R (Benway et al., 2023), WIDAR (Zhang et al., 2022), ExtraSensory (Gong et al., 2019; Hao et al., 2022), Stress-sensing(Eren and Navruz, 2022). Details are provided in Appendix A.1.
@@ 6.2. Metrics for evaluation
In summary, the metric $\Delta_{P}$ suggests how well the model $\mathcal{M}^{P^{a}_{i}}_{\theta}$ performs as compared to the generic model $\mathcal{M}^{G}_{\theta}$. Since the generic model has not learned the user-specific patterns. This metric quantifies CRoP’s ability to learn user-specific patterns. On the other hand, conventionally finetuned models $\mathcal{M}^{C^{a}_{i}}_{\theta}$ may learn user-specific patterns and forget the generic information of different contexts. Thus, the metric $\Delta_{G}$ quantifies CRoP’s ability to retain generic information.
@@ 6.3. Comparison with State-of-the-art
To demonstrate the efficacy of CRoP in achieving personalization $\Delta_{P}$ while maintaining generalization $\Delta_{G}$, we compare CRoP with 5 state-of-the-art approaches SHOT (Liang et al., 2020), PackNet (Mallya and Lazebnik, 2018), Piggyback (Mallya et al., 2018), CoTTA (Wang et al., 2022), and PTN (Burns et al., 2022).
@@
Figure 5. Detailed results for Stress sensing dataset User 3 (Scenario 1) through F1 score
@@ 6.3.3. Inferior generic model performance leads to higher gain in personalization
For ExtraSensory and Stress-sensing datasets, certain users experience sub-optimal performance with the generic model $\mathcal{M}^{G}_{\theta}$ and the personalized finetuning helped not only in available context but also in the unseen contexts. For instance, the generic model showed suboptimal performance for certain users such as user ‘80’ for ExtraSensory dataset under Scenario 1 and user ‘3’ for Stress-sensing dataset. Such users show higher benefit even in the unseen context when personalized using available context data, highlighting the importance of user-specific patterns.
@@ 7. Empirical Justification for CRoP
Model State User 0 User 1 User 2
$\mathcal{M}^{G}_{\theta}$ Generic 25 -389 12416
$\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}}$ Initial finetuned model (Alg1. Line 2) -973 - 1204 -5131
$\mathcal{M}^{P^{a}_{i}}_{\theta^{\downarrow}}$ Pruned (Alg1. Line 3) 3777 7002 -4197
$\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime\prime}}$ Mixed model (Alg1. Line 4) 4534 7254 8610
$\mathcal{M}^{P^{a}_{i}}_{\theta}$ Final model (Alg1. Line 5) 4534 7254 -2541
Generalizability introduced by the model mixing (Alg1. Line 5) step
$\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}}-\mathcal{M}^{P^{a}_{i}}_{\theta^{\downarrow}}$ Complementary to pruned model -285 -95 -2409
Corresponding Parameters in generic model 160 53 10297
Table 4. GIP for different model states for 3 users chosen for personalization for WIDAR Dataset.
Model $\mathcal{M}^{G}_{\theta}$ $\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}}$ $\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime\prime}}$ $\mathcal{A}(\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime\prime}},\mathcal{C})-\mathcal{A}(\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}},\mathcal{C})$ $\mathcal{M}^{P^{a}_{i}}_{\theta}$ $\mathcal{A}(\mathcal{M}^{P^{a}_{i}}_{\theta},\mathcal{C})-\mathcal{A}(\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}},\mathcal{C})$ $\mathcal{A}(\mathcal{M}^{P^{a}_{i}}_{\theta},\mathcal{C})-\mathcal{A}(\mathcal{M}^{G}_{\theta},\mathcal{C})$ Pruning
Generic Model Initial finetuning Mixed model Effect of mixing Final Model Final Model vs Initial Finetuning Final Model vs Generic Model amount
User $\mathcal{C}_{a}$ $\mathcal{C}_{u}$ $\mathcal{C}_{a}$ $\mathcal{C}_{u}$ $\mathcal{C}_{a}$ $\mathcal{C}_{u}$ $\mathcal{C}_{a}$ $\mathcal{C}_{u}$ $\mathcal{C}_{a}$ $\mathcal{C}_{u}$ $\mathcal{C}_{a}$ $\mathcal{C}_{u}$ $\mathcal{C}_{a}$ $\mathcal{C}_{u}$age
0 64.44 76.84 82.44 68.44 84.02 70.14 +1.58 +1.70 84.02 70.14 +1.58 +1.70 +19.58 -6.7 40
1 61.61 77.79 86.67 47.77 86.49 54.45 -0.18 +6.68 86.49 54.45 -0.18 +6.68 +24.88 -23.34 50
2 43.36 79.88 71.28 58.20 70.70 63.04 -0.58 +4.84 77.14 59.37 +5.86 +1.17 +33.78 -20.51 85
Average +0.27 +4.07 +2.42 +3.18 +26.08 -16.85
Table 5. Performance Comparison of model states after initial finetuning (Algorithm 1 line 2) $\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}}$ , model mixing (Algorithm 1 line 4) $\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime\prime}}$ and final finetuning (Algorithm 1 line 5) $\mathcal{M}^{P^{a}_{i}}_{\theta}$ for WIDAR dataset under Scenario 1.Effects of model mixing are computed by comparing Mixed model with initial finetuned model.
Euclidean Distance User0 User1 User2
Initial Finetuned Pruned Mixed Final Initial Finetuned Pruned Mixed Final Initial Finetuned Pruned Mixed Final
Generic model 10.05 13.09 1.62 1.62 10.31 17.73 2.16 2.16 19.85 61.26 4.36 15.18
Initial Finetuned model - 2.14 8.92 8.92 - 5.50 8.84 8.84 - 15.76 16.19 4.60
Pruned Model - - 11.47 11.47 - - 15.57 15.57 - - 56.91 31.46
Mixed Model - - - 0.0 - - - 0.0 - - - 6.48
Table 6. Euclidean distance between different model states for three users chosen for personalization for WIDAR dataset
@@ 7.1. How different steps of CRoP facilitate generalizability
This section empirically discusses how each step of CRoP (Algorithm 1) facilitates intra-user generalizability, enhancing similarity in the model’s behavior towards available $\mathcal{C}_{a}$ (available during personalization finetuning) and unseen contexts $\mathcal{C}_{u}$. The following sections outline the metrics, evaluation setup, and detailed analysis.
Metric for generalizability:
Shi et al. (2021) introduced the use of gradient inner product (GIP) to estimate the similarity between a model’s behavior across different domains. If the model incurs gradients $G_{i}$ and $G_{j}$ for two domains $D_{i}$ and $D_{j}$ respectively, the GIP $=G_{i}*G_{j}$ is estimated as $(\|G_{i}+G_{j}\|_{2})^{2}-(\|G_{i}\|_{2}^{2}+\|G_{j}\|_{2}^{2})$, where .$\|.\|_{2}$ signifies $\ell_{2}$-norm. The gradients $G_{i}$ and $G_{j}$ for the same model are computed using the general cross-entropy loss with a single backward pass on the samples corresponding to $D_{i}$ and $D_{j}$. The dimension of resulting gradients is the same as that of the model state. Through $\ell_{2}$-norm, the GIP computation considers gradients accumulated over all the parameters.
The sign of GIP, i.e., $G_{i}*G_{j}$, indicates how the model treats the two domains:
  • $G_{i}*G_{j}>0$ : signifies that the gradient for both domains has the same direction, which means that if a model takes a gradient step in the direction of $G_{i}$ or $G_{j}$, its performance across both domains $D_{i}$ and $D_{j}$ will improve. In essence, it reflects a higher generalizability.
  • $G_{i}*G_{j}<0$ : suggests the opposite directions of the gradients; thus, improving the model’s performance of one domain will worsen the performance of the other. This reflects lower generalizability.
Shi et al. (2021) also showed that higher GIP values with positive sign in two domains indicate more consistent behavior between the domains. Thus, we use GIP to quantify intra-user generalizability. Specifically, we treat the contexts $\mathcal{C}_{a}$ and $\mathcal{C}_{u}$ as two distinct domains, $D_{a}$ and $D_{u}$, and compute the gradients $G_{a}$ and $G_{u}$ for a given model state to calculate the GIP value.
Evaluation setup
This section’s evaluations consider the three users chosen for personalization for the WIDAR dataset under Scenario 1 which constitutes two non-overlapping contexts - the available context $\mathcal{C}_{a}$: Room-1, Torso Orientation- 1,2,3 and the unseen context $\mathcal{C}_{u}$: Room-1, Torso Orientation - 1,2,3 (as shown in Table 2). GIP is computed considering these contexts as distinct domains. While the previous sections presented results averaged across multiple seeds, this section presents results from a single random seed ($369284$) to provide a clear comparison of the model states obtained at each step of the algorithm.
Empirical result discussion
Table 6 shows the GIP of the generic model and the different states of the personalized model achieved after each step of Algorithm 1. To further support the discussion, Table 6 shows the inference accuracy for contexts $\mathcal{C}_{a}$ and $\mathcal{C}_{u}$ of these different model states along with the pruning amount incurred during the ToleratedPrune step. Additionally, Table 6 shows the Euclidean distances between different model states for a user, highlighting the parameter changes introduced at each step of Algorithm 1. According to the Tables 6, 6, and 6 results, a detailed analysis of the impact of each algorithm step is provided below.
  • Initial Finetuning (Algorithm 1 Line 2): This step aims to learn user-specific traits using limited context $\mathcal{C}_{a}$ data. While the accuracy of the initial finetuned model for context $\mathcal{C}_{a}$ improves as compared to the generic model, it drops significantly for unseen context $\mathcal{C}_{u}$. This pattern is consistent across all users, as shown in Table 6. Thus, initial finetuning results in generalizability loss which is supported by the drop in GIP values as shown in Table 6. Notably, User 2 experiences a much larger drop in GIP compared to Users 0 and 1. Table 6 shows that the generic model performs worse on User 2’s $\mathcal{C}_{a}$ compared to the others, suggesting that User 2’s user-specific data patterns differ significantly from the generic patterns. Thus, User 2 requires greater parameter adjustments (i.e., larger gradient updates) as evidenced by a higher Euclidean distance of $19.85$ units between the generic and finetuned model as compared to approximately $10$ units for others (Table 6), leading to a higher difference in GIP between the generic and finetuned models.
  • Tolerated Prune (Algorithm 1 Line 3): This module aims to find a maximal pruning amount for the personalized model without losing performance in the context $\mathcal{C}_{a}$. These pruning amounts vary among different users. For instance, User 0 and 1 undergo $40\%$ and $50\%$ pruning, respectively, while User 2 undergoes an extensive pruning amount of $85\%$ (shown in Table 6). This is so because the training loss contains a regularization component (Algorithm 1 Line 2), User 2 undergoes higher parameter adjustments, i.e., larger gradient updates, which in turn results in higher regularization. Thus, the important weights learned for User 2 are restricted to a smaller subnetwork, which contributes to approximately $15\%$ parameters, thus allowing higher pruning. This is further supported by the higher Euclidean distance between the initial-finetuned and pruned model for User 2 as compared to other users (Table 6).
    The effect of the regularizer on the variable pruning amount highlights a key aspect of Algorithm 1. For individuals requiring larger parameter adjustments (through larger gradient updates) —indicated by significant deviations of model parameters from the generic model—the risk of losing generalizability increases. In such cases, the regularizer ensures higher polarization of parameters allowing higher pruning, restoring more parameters from the generic model during the subsequent mixing steps, which we will demonstrate for User 2 in the following discussion. This facilitates greater recovery, enhancing generalizability.
    Additionally, model pruning has been shown to enhance generalization in literature(Jin et al., 2022), and we also observed an increase in GIP value in the pruned model (Step $3$) as shown in Table 6. On further analyzing the sub-network of parameters selected for removal during the pruning step ($\mathcal{M}^{P^{a}_{i}}_{\theta^{\prime}}-\mathcal{M}^{P^{a}_{i}}_{\theta^{\downarrow}}$, complementary to the pruned model), we observed that the GIP corresponding to this subnetwork was negative (in Table 6), contributing to discrepancy in model behavior, and removal of these parameters improves generalizability.
  • Model Mixing (Algorithm 1 Line 4): In this step, the parameters removed during the previous step are replaced by corresponding parameters from the generic model. Table 6 shows that this step further improves GIP. The details for this increase are discussed below.
    The parameters removed in the previous step had negative GIP, while the corresponding parameters in the generic model alone exhibited positive GIP (Table 6, last row). Restoring these parameter values from the generic model enhances GIP in the mixed models, as shown in Table 6, improving the generalizability. Additionally, the reduced distance between the generic and mixed models (Table 6) indicates that this step aligns the personalized model closer to the more generalizable generic model. Notably, for User 2, the distance reduction is significantly greater due to a larger number of parameters restored from the generic model. This restoration leads to a substantial increase in the GIP value (Table 6), clearly demonstrating improved generalization.
  • Final Finetuning (Algorithm 1 Line $5$): Finally, as discussed in Section 5.2, model mixing may result in inconsistencies in the model parameters’ signs and gradients which can have a degrading impact on model performance for some users. We employ the final finetuning to address that. As evident from table 6, User 0 and User 1 do not require fine-tuning, as indicated by the Euclidean distance between the final and the mixed model remaining 00. However, User 2 does undergo final fine-tuning, as shown by a non-zero difference between the mixed and final model in Table 6, further improving the overall performance.
    Final fine-tuning may slightly reduce generalizability compared to the mixed model. Table 6 supports this, showing a drop in GIP for the final fine-tuned model, leading to lower performance in $\mathcal{C}_{u}$ for User 2. However, overall performance across $\mathcal{C}_{a}$ and $\mathcal{C}_{u}$ remains higher than the mixed model. More importantly, despite this drop, the final model outperforms the initial fine-tuned model (from Line 2 in Algorithm 1) in both GIP and performance on $\mathcal{C}_{a}$ and $\mathcal{C}_{u}$, indicating improved generalization and personalization.
7.2. TSNE visualization
This section further demonstrates the generalizability benefits of different steps of Algorithm 1 by employing t-SNE plots to visualize feature separability across classes in different model states. Since t-SNE represents higher dimensional features to a 2-dimensional space, we select the ExtraSensory dataset, with is a binary classification problem (walking vs sitting activity) for ease of observability. Figure 6 presents t-SNE plots for four model states—generic, initial fine-tuned, mixed, and final—showing features along with their class membership (indicated by colors). These features are extracted for both available and unseen contexts for user ‘7c,’ personalized using available context data $\mathcal{C}_{a}$ under Scenario 2, and derived from a GRU-based model trained on the ExtraSensory dataset. The model constitutes two parts: feature-extractor and classifier. The feature-extractor contains three batch-normalized 1D convolution layers followed by a linear layer that feeds into a batch- normalized recursive (GRU) layer and two linear layers to generate embeddings. The embeddings generated by the feature extractor are used to generate the t-SNE plots. (model details are provided in Appendix A.1).
Figure 6(a) shows that the features of different classes extracted by the generic model for the available context $\mathcal{C}_{a}$ have high overlap, leading to a low accuracy of approximately $73\%$. On the other hand, the same generic model shows better separability for unseen context $\mathcal{C}_{u}$, resulting in an accuracy of around $92\%$ (Figure 6(e)). As discussed in earlier sections, finetuning the generic model improves the model’s performance in the available context $\mathcal{C}_{a}$, which is evident from $6\%$ gain in accuracy in $\mathcal{C}_{a}$ and better separability of features as shown in Figure 6(b). At the same time, unseen context $\mathcal{C}_{u}$ suffers, resulting in a significant accuracy drop of around $25\%$ and an increase in the overlap between the extracted features for the two classes, as shown in Figure 6(f). Next, the model mixing step brings generic information from the generic model to the personalized model, and as expected, an improvement in the model’s performance for $\mathcal{C}_{u}$ can be seen through a gain in accuracy and separability of features (Figure 6(g)). Simultaneously, as discussed in Section 5.2, the inconsistencies in the mixed model result in performance degradation, especially for context $\mathcal{C}_{a}$ (Figure 6(c)). Final finetuning aims to address this issue, and as depicted by Figure 6(d), a significant performance gain due to enhancement of separability for context $\mathcal{C}_{a}$ can be obtained without harming the model’s performance (and separability of features among different classes) for the unseen context $\mathcal{C}_{u}$ as shown in 6(h).
Thus, the level of feature separability across $\mathcal{C}_{a}$ and $\mathcal{C}_{u}$ for different model states reflects the generalizability improvements achieved through the steps of Algorithm 1.
(a) $\mathcal{C}_{a}$ Generic Model
Accuracy = $72.99\%$
(b) $\mathcal{C}_{a}$ Initial finetuned
Accuracy = $78.61\%$
(c) $\mathcal{C}_{a}$- Mixed Model
Accuracy = $68.74\%$
(d) $\mathcal{C}_{a}$ Final Model
Accuracy = $83.18\%$
(e) $\mathcal{C}_{u}$ Generic Model
Accuracy = $92.32\%$
(f) $\mathcal{C}_{u}$ Initial finetuned
Accuracy = $67.26\%$
(g) $\mathcal{C}_{u}$ Mixed Model
Accuracy = $88.62\%$
(h) $\mathcal{C}_{u}$ Final Model
Accuracy = $89.39\%$
Figure 6. t-SNE plots for user ’7c’ in ExtraSensory Dataset for Scenario 2 where avalable context $\mathcal{C}_{a}$ (‘phone in pocket’ and ’phone in bag’) and $\mathcal{C}_{u}$ (’phone in hand’)context. The colors of the samples indicate their class membership.